I dont know what section I need, but can someone help me with the following problem? I have no clue where to even begin. Thank you.
Find f(x+h)-f(x)/h for the following function:
3x-2

Answer by stanbon(75887)

Find f(x+h)-f(x)/h for the following function:
f(x) = 3x-2
f(x+h) = 3(x+h)-2 = 3x + 3h -2
-----
f(x+h)-f(x) = 3h
-----------------------
[f(x+h)-f(x)]/h = 3h/h = 3
============================================
Cheers,
Stan H.